OpenStudy (anonymous):

A motorcyclist changes the velocity of his bike from 20.0 meters/second to 35.0 meters/second under a constant acceleration of 4.00 meters/second^2. How long does it take the bike to reach the final velocity? a) 2.02 seconds b) -2.02 seconds c) 3.75 seconds d) -3.75 seconds

OpenStudy (anonymous):

\[ v_0=20{\rm m/s}\\ v_f=35{\rm m/s}\\ a=4{\rm m/s^2}\\ t=? \] from the definition of acceleration, \[a=\frac{v_f-v_0}{t}\]

OpenStudy (anonymous):

@graceface are u satisfied with @electrokid answer???

OpenStudy (anonymous):

I'm still trying to figure it out lol.